# Nearly optimal spectral gaps for random Belyi surfaces

### v1

@@ Abstract.
In this paper, we show that for any $\epsilon>0$, a random hyperbolic surface in the Brooks-Makover model has a spectral gap greater than $\frac{1}{4}-\epsilon$, confirming the nearly optimal spectral gap conjecture in this model.
@@ 1. Introduction
Weil-Petersson Model: The study of spectral gaps for random closed hyperbolic surfaces in this model began with Mirzakhani [Mir13], who proved a lower bound $0.00247$. This bound was significantly improved to $\left(\frac{3}{16}-\epsilon\right)$ for any $\epsilon>0$ by Wu-Xue [WX22] using a completely different method, a result independently obtained by Lipnowski-Wright [LW24]. In a series of subsequent works, Anantharaman-Monk raised the lower bound further to $\left(\frac{2}{9}-\epsilon\right)$ in [AM23] and eventually to $\left(\frac{1}{4}-\epsilon\right)$ in the remarkable work [AM25], resolving the nearly optimal spectral gap conjecture for the Weil-Petersson model. Most recently, Hide-Macera-Thomas [HMT25] provided an alternative proof via the polynomial method, additionally refining the $\epsilon$ term to a significant explicit polynomial error of order $\frac{1}{g^{\delta}}$ for some uniform $\delta>0$.
Covering Model: Magee-Naud-Puder [MNP22] proved that for any $\epsilon>0$ and any fixed closed hyperbolic surface $X$, the eigenvalues (with multiplicity) of a random covering surface of $X$ within $\left(0,\frac{3}{16}-\epsilon\right)$ coincide with those of $X$. In a recent remarkable work [MPvH25], Magee-Puder-van Handel used the polynomial method to improve the bound from $\left(\frac{3}{16}-\epsilon\right)$ to $\left(\frac{1}{4}-\epsilon\right)$, resolving the nearly optimal spectral gap conjecture for the covering model.
@@
For any $\epsilon>0$,
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \lambda_{1}\left(S^{C}(\Gamma,\mathcal{O})\right)>\frac{1}{4}-\epsilon\right)=1.$
It is known from [BM04] that a random cusped hyperbolic surface $S^{O}(\Gamma,\mathcal{O})$ has the so-called large cusps property. This together with [BM01] tells that as $n\to\infty$, for a random $(\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star}$,
$\lambda_{1}\left(S^{C}(\Gamma,\mathcal{O})\right)\geq\min\left\{\frac{1}{4}-\epsilon,(1-o(1))\cdot\lambda_{1}\left(S^{O}(\Gamma,\mathcal{O})\right)\right\}.$
Therefore, to prove Theorem 1.1, it suffices to show that a random cusped hyperbolic surface $S^{O}(\Gamma,\mathcal{O})$ has no eigenvalues in $\left(0,\frac{1}{4}-\epsilon\right)$ for any $\epsilon>0$. Our approach is inspired by the work [MPvH25] of Magee-Puder-van Handel that builds upon the polynomial method in e.g. [CGVTvH24, CGVvH24, MdlS24]. Meanwhile, we believe that the novel ingredients developed in this work will find further applications in studying the asymptotics of other geometric quantities in this model of random hyperbolic surfaces.
For more related results of spectral gaps on random hyperbolic surfaces, one may refer to e.g. [Hid23, HT25, SW25] for the Weil-Petersson model and to e.g. [MN20, HM23, LM25] for the covering model. Similar results on spectral gaps also hold for random graphs. Alon [Alo86] conjectured and Friedman [Fri08] proved that a random $d$-regular $(d\geq 3)$ graph has nearly optimal spectral gap, as the number $N$ of vertices goes to infinity. A recent breakthrough of Huang-Mckenzie-Yau [HMY25] proved that as $N\to\infty$, about $69\%$ $d$-regular graphs with $N$ vertices are Ramanujan graphs, that is, have optimal spectral gaps. It is natural to ask whether an analogous result holds for all these three models of random closed hyperbolic surfaces. That is, does there exist a constant $c\in(0,1)$ such that
@@
where the constant $c>0$ may be different for these three models of random closed hyperbolic surfaces. The existence of infinite families of Ramanujan graphs was first proved in [LPS88, Mar88, MSS15]. We remark here that it remains even open that there exists a closed hyperbolic surface $X_{g}$ of genus $g$ such that $\lambda_{1}(X_{g})\geq\frac{1}{4}$ for all $g\geq 2$.
@@
$\textnormal{tr }\Pi_{\gamma}(\sigma,\tau):=\frac{1}{6n}\textnormal{Tr }\left(\textnormal{std}\circ\Phi(\sigma,\tau)(\gamma)\right),$
@@
There exist a universal constant $C\geq 1$ and a sequence $\{u_{i}(\gamma)\}_{i\geq 0}\subset\mathbb{R}$ for all $\gamma\in\textnormal{PSL}(2,\mathbb{Z})$ such that $u_{0}(\gamma)=\rho(\lambda(\gamma))$ and
$\left|\mathbb{E}_{n}\left[\textnormal{tr }\Pi_{\gamma}\right]-\sum\limits_{i=0}^{m-1}\frac{u_{i}(\gamma)}{(6n)^{i}}\right|\leq\frac{(Cm)^{Cm}}{(6n)^{m}}$
for all $m\geq|\gamma|+1$ and $n\geq Cm^{C}$.
@@
It is known from e.g. [Ber11, Theorem 3.38] that $\lambda_{1}(\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z}))>\frac{1}{4}$. Finally, based on Proposition 1.2 and 1.3, we finish the proof of Theorem 1.1 following the arguments in [MPvH25, HM23, HTr05].
@@
Acknowledgment. We are very grateful to Bram Petri for invaluable discussions on the new description of the Brooks-Makover model. We also thank Yuxin He, Yuhao Xue and Haohao Zhang for many helpful discussions and suggestions on this project. The first named author is supported by the NSFC grant No. 12401081, and the second named author is partially supported by the NSFC grant No. 12425107 and 12361141813.
@@ 2. Brooks-Makover Model
In this section, we recall the construction of random hyperbolic surfaces by Brooks-Makover in [BM04], and introduce an alternative description in terms of the covering surfaces.
@@ 2.1. Brooks-Makover random hyperbolic surfaces
2.2. Large cusps condition
Given $(\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star}$, we let $\{L_{i}\}_{i\in\mathcal{I}(\Gamma,\mathcal{O})}$ denote the set of all canonical horocycle loops of $S^{O}(\Gamma,\mathcal{O})$. Write the index set $\mathcal{I}(\Gamma,\mathcal{O})$ as $\mathcal{I}$ for simplicity. For each $i\in\mathcal{I}$, the canonical horocycle loop $L_{i}$ bounds a punctured disk $N_{i}\subset S^{O}(\Gamma,\mathcal{O})$ with a puncture $p_{i}$. Moreover $D_{i}=N_{i}\cup\{p_{i}\}$ is an open topological disk in $S^{C}(\Gamma,\mathcal{O})$.
Now we recall the so-called large cusps condition defined in [Bro99, BM04]. For any $i\in\mathcal{I}$, up to a conjugation one can lift the puncture $p_{i}$ to $\infty$ in the boundary $\partial\mathbb{H}$ of the hyperbolic upper half-plane $\mathbb{H}$. Let $\pi_{i}:\mathbb{H}\to S^{O}(\Gamma,\mathcal{O})$ be the covering map with $\pi_{i}(\infty)=p_{i}$ and assume that the Mobius transformation through doing one turn around $p_{i}$ corresponds to the parabolic isometry $z\mapsto z+1$. For any $l>0$, set
$C_{i}(l)=\pi_{i}\left(\left\{z\in\mathbb{H};\ \textnormal{Im}\ z\geq\frac{1}{l}\right\}\right)$
and
$h_{i}(l)=\pi_{i}\left(\left\{z\in\mathbb{H};\ \textnormal{Im}\ z=\frac{1}{l}\right\}\right).$
Then for suitable $l>0$, the set $C_{i}(l)$ is a cusp around $p_{i}$ whose boundary curve $h_{i}(l)$ is the projection of a horocycle. We also denote by $C^{l}$ the standard cusp
$C^{l}=\left\{z\in\mathbb{H};\ \textnormal{Im }z\geq\frac{1}{l}\right\}/(z\sim z+1)$
endowed with the hyperbolic metric. Now we recall [Bro99, Definition 2.1] or [BM04, Definition 3.1] saying that.
Definition.
Given any $l>0$, a hyperbolic surface $S^{O}$ has cusps of length $\geq l$ if
  1. (i)
    $C_{i}(l)$ is isometric to $C^{l}$ for any $i\in\mathcal{I}$;
  2. (ii)
    $C_{i_{1}}(l)\cap C_{i_{2}}(l)=\emptyset$ for any $i_{1}\neq i_{2}\in\mathcal{I}$.
Brooks and Makover in [BM04] proved that for any $l>0$, as $n\to\infty$ a generic hyperbolic surface $S^{O}(\Gamma,\mathcal{O})$ has cusps of length $\geq l$. More precisely,
@@ 2.2. Large cusps condition
(([BM04, Theorem 2.1A])) Let $(\Gamma,\mathcal{O})$ be a random element of $\mathcal{F}^{\star}_{n}$. Then for any $l>0$,
$\lim\limits_{n\to\infty}\textnormal{Prob}^{n}_{\textnormal{BM}}\left\{S^{O}(\Gamma,\mathcal{O})\text{ has cusps of length}\geq l\right\}=1.$
Assume that $S^{O}$ is a non-compact hyperbolic surface and $S^{C}$ is the conformal compactification of $S^{O}$. If $S^{O}$ has large cusps, then the geometries of $S^{O}$ and $S^{C}$ are very close in their thick parts. Concerning eigenvalues, we have the following lemma.
Lemma 2.2.
(([BM01, Lemma 1.2])) For all $l$ sufficient large and for all $\epsilon>0$, there there exists a constant $C(l,\epsilon)>0$ with the following property: if $S^{O}$ has cusps of length $\geq l$, then
$\lambda_{1}\left(S^{C}\right)\geq\min\left\{\frac{1}{4}-\epsilon,\ C(l,\epsilon)\lambda_{1}\left(S^{O}\right)\right\}.$
Furthermore $C(l,\epsilon)\to 1$ as $l\to\infty$.

### v2

@@ Abstract.
In this paper, we show that a random hyperbolic surface in the Brooks-Makover model has a spectral gap greater than $\left(\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)$, confirming the nearly optimal spectral gap conjecture in this model.
@@ 1. Introduction
Weil-Petersson Model: The study of spectral gaps for random closed hyperbolic surfaces in this model began with Mirzakhani [Mir13], who proved a lower bound $0.00247$. This bound was significantly improved to $\left(\frac{3}{16}-\epsilon\right)$ for any $\epsilon>0$ by Wu-Xue [WX22] using a completely different method, a result independently obtained by Lipnowski-Wright [LW24]. In a series of subsequent works, Anantharaman-Monk raised the lower bound further to $\left(\frac{2}{9}-\epsilon\right)$ in [AM23] and eventually to $\left(\frac{1}{4}-\epsilon\right)$ in the remarkable work [AM25], resolving the nearly optimal spectral gap conjecture for the Weil-Petersson model. Most recently, Hide-Macera-Thomas [HMT25b] provided an alternative proof via the polynomial method, additionally refining the $\epsilon$ term to a significant explicit polynomial error of order $\frac{1}{g^{\delta}}$ for some uniform $\delta>0$.
Covering Model: Magee-Naud-Puder [MNP22] proved that for any $\epsilon>0$ and any fixed closed hyperbolic surface $X$, the eigenvalues (with multiplicity) of a random covering surface of $X$ within $\left(0,\frac{3}{16}-\epsilon\right)$ coincide with those of $X$. In a recent remarkable work [MPvH25], Magee-Puder-van Handel used the polynomial method to improve the bound from $\left(\frac{3}{16}-\epsilon\right)$ to $\left(\frac{1}{4}-\epsilon\right)$, resolving the nearly optimal spectral gap conjecture for the covering model. Hide-Macera-Thomas [HMT25a] refined the term $\epsilon$ to a polynomial error of order $\frac{1}{g^{b}}$ for some uniform $b>0$.
@@
The following limit holds:
$\lim\limits_{n\to\infty}\textnormal{Prob}_{\textnormal{BM}}^{n}\left((\Gamma,\mathcal{O})\in\mathcal{F}_{n}^{\star};\ \lambda_{1}\left(S^{C}(\Gamma,\mathcal{O})\right)>\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)=1.$
We will first show that the spectral gap of a cusped hyperbolic surface becomes larger after compactification (see Proposition 2.1). Therefore, to prove Theorem 1.1, it suffices to show that a random cusped hyperbolic surface $S^{O}(\Gamma,\mathcal{O})$ has no eigenvalue in $\left(0,\frac{1}{4}-\left(\frac{1}{n}\right)^{\frac{1}{221}}\right)$. Our approach is inspired by the work [MPvH25] of Magee-Puder-van Handel that builds upon the polynomial method in e.g. [CGVTvH24, CGVvH24, MdlS24]. Meanwhile, we believe that the novel ingredients developed in this work will find further applications in studying the asymptotics of other geometric quantities in this model of random hyperbolic surfaces. The power $\frac{1}{221}$ is far from optimal. It would be interesting to study the optimal decay rate.
For more related results of spectral gaps on random hyperbolic surfaces, one may refer to e.g. [Hid23, HT25, SW25] for the Weil-Petersson model and to e.g. [MN20, HM23, LM25] for the covering model. Similar results on spectral gaps also hold for random graphs. Alon [Alo86] conjectured and Friedman [Fri08] proved that a random $d$-regular $(d\geq 3)$ graph has nearly optimal spectral gap, as the number $N$ of vertices goes to infinity. A recent breakthrough of Huang-Mckenzie-Yau [HMY25] proved that as $N\to\infty$, about $69\%$ $d$-regular graphs with $N$ vertices are Ramanujan graphs, that is, have optimal spectral gaps. It is natural to ask whether an analogous result holds for all these three models of random closed hyperbolic surfaces. That is,
Question 1.
Does there exist a constant $c\in(0,1)$ such that
@@
where the constant $c>0$ may be different for these three models of random closed hyperbolic surfaces.
The existence of infinite families of Ramanujan graphs was first proved in [LPS88, Mar88, MSS15]. We remark here that it remains even open that there exists a closed hyperbolic surface $X_{g}$ of genus $g$ such that $\lambda_{1}(X_{g})\geq\frac{1}{4}$ for all $g\geq 2$. It is known from [SU13, Coo18, FBP23] that such so-called Ramanujan surfaces exist in genera $2$ to $7$, $14$ and $17$. In this paper, based on [BLS20] we show that Ramanujan surfaces also exist in genera
$g(N)=1+\frac{N^{2}(N-6)}{24}\prod_{p\mid N}\Big(1-\frac{1}{p^{2}}\Big)$
whenever $7\leq N\leq 226$. See Theorem 2.5.
@@
$\textnormal{tr }\Pi_{\gamma}(\sigma,\tau)\overset{\mathrm{def}}{=}\frac{1}{6n}\textnormal{Tr }\left(\textnormal{std}\circ\Phi(\sigma,\tau)(\gamma)\right),$
@@
There exists a sequence $\{u_{i}(\gamma)\}_{i\geq 0}\subset\mathbb{R}$ for all $\gamma\in\textnormal{PSL}(2,\mathbb{Z})$ such that $u_{0}(\gamma)=\rho(\lambda(\gamma))$ and
$\left|\mathbb{E}_{n}\left[\textnormal{tr }\Pi_{\gamma}\right]-\sum\limits_{i=0}^{m-1}\frac{u_{i}(\gamma)}{(6n)^{i}}\right|\leq\frac{(18m)^{18m}}{(6n)^{m}}$
for all $m\geq|\gamma|+1$ and $n\geq 18m^{18}$.
@@
It is known from e.g. [Ber11, Theorem 3.38] that $\lambda_{1}(\mathbb{H}/\textnormal{PSL}(2,\mathbb{Z}))>\frac{1}{4}$. Finally, based on Proposition 1.2 and 1.3, we finish the proof of Theorem 1.1 following the arguments in [HMT25a, MPvH25].
@@
Acknowledgment. We are very grateful to Bram Petri for invaluable discussions on the new description of the Brooks-Makover model and his comment on this paper. We also thank Yuxin He, Yuhao Xue and Haohao Zhang for many helpful discussions and suggestions on this project. We are also grateful to Doron Puder, Joe Thomas and Ramon van Handel for their interests and comments on this paper. The first named author is supported by the NSFC grant No. 12401081, and the second named author is partially supported by the NSFC grant No. 12425107 and 12361141813.
@@ 2. Brooks-Makover Model
In this section, we recall the construction of random hyperbolic surfaces by Brooks-Makover in [BM04], and introduce an alternative description.
@@ 2.1. Brooks-Makover random hyperbolic surfaces
2.2. Comparison of eigenvalues
In this subsection, we provide the following result that says that the spectral gap becomes larger after conformal compactification. More precisely,
@@ 2.2. Comparison of eigenvalues
Let $S^{O}$ be a non-compact hyperbolic surface of finite area, and let $S^{C}$ be the conformal compactification of $S^{O}$ obtained by filling in all the cusps of $S^{O}$, then
$\lambda_{1}(S^{C})>\textnormal{RayQ}(S^{O})\overset{\mathrm{def}}{=}\inf\limits_{f\in L^{2}(S^{O}),\ \int_{S^{O}}f=0}\frac{\int_{S^{O}}|\nabla f|^{2}\mathop{\rm dvol}_{O}}{\int_{S^{O}}f^{2}\mathop{\rm dvol}_{O}},$
where $\mathop{\rm dvol}_{O}$ is the volume form on $S^{O}$.
Proof.
Denote by $\mathop{\rm dvol}_{C}$ the volume form on $S^{C}$. There is a natural holomorphic inclusion
$(S^{O},\mathrm{dvol_{O}})\to(S^{C},\mathrm{dvol_{C}}).$
By the Ahlfors–Schwarz Lemma,
$\mathrm{dvol_{O}}>\mathrm{dvol_{C}}.$
Let $\varphi\in C^{\infty}(S^{C})$ be an eigenfunction w.r.t the first eigenvalue $\lambda_{1}(S^{C})$. Define
$\widetilde{\varphi}\overset{\mathrm{def}}{=}\varphi-\frac{1}{\mathrm{vol}(S^{O})}\int_{S^{O}}\varphi\mathrm{dvol_{O}}.$
Then we have $\int_{S^{O}}\widetilde{\varphi}\mathrm{dvol_{O}}=0$. Moreover, we have
(2) $\int_{S^{O}}(\widetilde{\varphi})^{2}\mathrm{dvol_{O}}$ $>\int_{S^{C}}(\widetilde{\varphi})^{2}\mathrm{dvol_{C}}$
$=\int_{S^{C}}\varphi^{2}\mathrm{dvol_{C}}+\frac{\mathrm{vol}(S^{C})}{\mathrm{vol}(S^{O})^{2}}\Big(\int_{S^{O}}\varphi\mathrm{dvol_{O}}\Big)^{2}$
$\geq\int_{S^{C}}\varphi^{2}\mathrm{dvol_{C}},$
where in the second line we use $\int_{S^{C}}\varphi\mathrm{dvol_{C}}=0$. By the conformal invariance of two-dimensional Dirichlet integral, we have
(3) $\int_{S^{O}}\lvert\nabla\widetilde{\varphi}\rvert^{2}\mathrm{dvol_{O}}=\int_{S^{C}}\lvert\nabla\widetilde{\varphi}\rvert^{2}\mathrm{dvol_{C}}=\int_{S^{C}}\lvert\nabla\varphi\rvert^{2}\mathrm{dvol_{C}}.$
It then follows from (2) and (3) that
$\lambda_{1}(S^{C})=\frac{\int_{S^{C}}\lvert\nabla\varphi\rvert^{2}\mathrm{dvol_{C}}}{\int_{S^{C}}\varphi^{2}\mathrm{dvol_{C}}}>\frac{\int_{S^{O}}\lvert\nabla\widetilde{\varphi}\rvert^{2}\mathrm{dvol_{O}}}{\int_{S^{O}}(\widetilde{\varphi})^{2}\mathrm{dvol_{O}}}\geq\textnormal{RayQ}(S^{O}).$
The proof is complete. ∎
The following fundamental question still remains open (see e.g. [Sar03]).
Question 2.
Does a complete non-compact hyperbolic surface of finite area always have a non-zero eigenvalue?
Combined with the standard result for spectral theory (see e.g. [RS78, Theorem XIII.1]), we have the following consequence of Proposition 2.1 that gives a new insight into Question 2.
Theorem 2.2.
Let $S^{O}$ be a non-compact hyperbolic surface of finite area such that its conformal compactification $S^{C}$ has first eigenvalue $\leq\frac{1}{4}$. Then $S^{O}$ has a non-zero eigenvalue.
Following the terminology of [FBP23], we call a finite-area hyperbolic surface $X$ a Ramanujan surface if $\lambda_{1}(X)\geq\frac{1}{4}$. As introduced in the introduction, the following question is important and widely open (see e.g. [FBP23, van25]).
Question 3.
Does there exist a closed Ramanujan surface of genus $g$ for all $g\geq 2$?
Such surfaces are known to exist in genera $2$, 33, and $4$, as established in [SU13] and [Coo18], where they made numerical computations of $\lambda_{1}$ for the Bolza surface, the Klein quartic and the Bring’s curve, respectively. More recently, in [FBP23], Bourque–Petri using linear programming methods proved the existence of closed Ramanujan surfaces in genera $5$ to $7$, $14$ and $17$.
Now we recall some arithmetics. The principal congruence subgroup $\Gamma(N)$ of the modular group $\mathrm{PSL}(2,\mathbb{Z})$ of level $N\geq 2$ is defined as
$\Gamma(N)\overset{\mathrm{def}}{=}\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{PSL}(2,\mathbb{Z});\ \begin{pmatrix}a&b\\ c&d\end{pmatrix}\equiv\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\mod N\right\}.$
The quotient $Y(N):=\mathbb{H}/\Gamma(N)$ is a non-compact hyperbolic surface of finite area. The genus $g(N)$ of $Y(N)$ can be computed explicitly by using the Riemann–Hurwitz formula, see e.g. [Shi71] or [DS05, Page 108]. For $N\geq 3$, an explicit formula is
$g(N)=1+\frac{N^{2}(N-6)}{24}\prod_{p\mid N}\Big(1-\frac{1}{p^{2}}\Big).$
This is always a non-negative integer.
The celebrated Selberg $\frac{1}{4}$-eigenvalue conjecture states that
Conjecture 2.3.
For each integer $N\geq 1$,
$\lambda_{1}(Y(N))\geq\frac{1}{4}.$
We refer the reader to [Sel65, GJ78, LRS95, KS02a, KS02b, Kim03] for the history and developments on this conjecture. Huxley [Hux85] proved the Selberg eigenvalue conjecture for $N\leq 18$. Recently, Booker-Lee-Strömbergsson [BLS20] greatly improved Huxley’s result.
Theorem 2.4 ([BLS20, Theorem 1.1]).
The Selberg eigenvalue conjecture holds for the principal congruence subgroups $\Gamma(N)$ with $N\leq 226$.
Indeed, they proved that the conjecture also holds for another family of congruence subgroups $\Gamma_{1}(N)$ with $N\leq 880$, improving the previous result of Booker–Strömbergsson [BS07]. The genera of the corresponding surfaces $\mathbb{H}/\Gamma_{1}(N)$ can be also explicitly computed (see e.g. [DS05, Page 108]); however, the resulting table for $N\leq 880$ is too long and is therefore omitted here.
Let $X(N)$ be the conformal compactification of $Y(N)$ obtained by filling all the cusps. Then by Proposition 2.1 and noting that $g(N)>1$ when $N\geq 7$, we have
Theorem 2.5.
For each $7\leq N\leq 226$, we have
$\lambda_{1}(X(N))>\frac{1}{4}.$
In particular, there exist closed Ramanujan surfaces in all the corresponding genera (see Table 1).
NN $g(N)$     NN $g(N)$     NN $g(N)$     NN $g(N)$     NN $g(N)$     NN $g(N)$
3 0     42 1729     81 18226     120 43777     159 143209     198 207361
4 0     43 2850     82 15961     121 69576     160 118273     199 318451
5 0     44 2281     83 22100     122 53941     161 163681     200 232801
6 1     45 2809     84 14977     123 65521     162 113725     201 291721
7 3     46 2641     85 22753     124 56641     163 173800     202 249901
8 5     47 3773     86 18481     125 74376     164 132721     203 330961
9 10     48 2689     87 22681     126 51841     165 152641     204 228097
10 13     49 4215     88 19681     127 81313     166 137761     205 334321
11 26     50 3301     89 27391     128 62465     167 187083     206 265201
12 25     51 4321     90 18145     129 75769     168 124417     207 318385
13 50     52 3865     91 28561     130 62497     169 192830     208 271489
14 49     53 5500     92 22705     131 89376     170 141697     209 365401
15 73     54 3889     93 27841     132 60481     171 178201     210 235009
16 81     55 5881     94 24289     133 91441     172 153385     211 380276
17 133     56 4801     95 32041     134 71809     173 208250     212 289225
18 109     57 6121     96 23041     135 83593     174 141121     213 347761
19 196     58 5461     97 35673     136 74881     175 202801     214 297649
20 169     59 7686     98 27049     137 102443     176 163201     215 386233
21 241     60 5185     99 33481     138 69697     177 198361     216 272161
22 241     61 8526     100 28201     139 107066     178 170281     217 405121
23 375     62 6721     101 40376     140 77185     179 230956     218 314821
24 289     63 8209     102 27649     141 99361     180 150337     219 378289
25 476     64 7425     103 42875     142 85681     181 238876     220 308161
26 421     65 9913     104 32929     143 115081     182 177409     221 433441
27 568     66 7201     105 38017     144 79489     183 219481     222 295489
28 529     67 11408     106 35101     145 116761     184 187969     223 449625
29 806     68 8929     107 48178     146 93241     185 244873     224 334849
30 577     69 11089     108 33049     147 110545     186 172801     225 394201
31 1001     70 9217     109 50986     148 97129     187 260641     226 351121
32 833     71 13651     110 37441     149 132276     188 200929
33 1081     72 9505     111 47881     150 86401     189 237169
34 1009     73 14875     112 40705     151 137751     190 198721
35 1393     74 11629     113 56925     152 105121     191 281201
36 1081     75 13801     114 38881     153 127009     192 190465
37 1768     76 12601     115 57553     154 106561     193 290225
38 1441     77 17041     116 46201     155 143041     194 221089
39 1849     78 12097     117 55945     156 100801     195 254017
40 1633     79 18981     118 48721     157 155078     196 223441
41 2451     80 14209     119 65089     158 118561     197 308848
Table 1. The values of $g(N)$ for $3\leq N\leq 226$